Reports 1-1 of 1 Clear search Modify search
AdV-SVS (Pre-commissioning)
garaventa, sorrentino - 1:01 Friday 10 September 2021 (53118) Print this report
EQB1 activity on: the overlapping of SC-BAB, the centering of them on HD_PDs and the BS splitting ratio balancing

During the shift of this morning (09/09/2021) we worked on:

  • the recovery of the beatnote signal on EQB1_HD_RF_DEMOD, moving the mirror in front of this PD (power_BAB about 2mW, power_SC about 1mW) and so engaging the SC_AA loop to improve this signal on EQB1_HD_RF_DEMOD (see Fig.1); in order to improve this signal, we also moved L2 of SC telescope (see Fig.2).

  • Then, we open SC_AA loop (process EQB1_HD_AA), we moved the IR shutter in order to send beams on EQB1_IR_PD_MONI, so we closed LO shutter, and we scanned the OPA_p_pol (process SQZ_CTRL) removing temporarily BAB. Fig.3 shows MM of about 97% (TEM00 reflection peak is about 170 mV, the other strongest HOM is about 5mV).

  • With LO shutter closed, we checked the BS spitting ratio and there is an imbalance of about 1%.

  • Then, we worked on the centering of BAB-SC on HD_PDs following this procedure:

  1. we sent fast modulation on M4X and slow modulation on M4Y;

  2. we moved HD_PD1,2 in their 3 DOFs 10 steps at a time in order to center the photodiodes (see Fig.4).

  3. on EQB1_HD_DIFF_RF_sample we saw the FFT in order to check the amplitude of the beatnote at 4MHz (Fig.5 without LO, Fig.6 with LO).

Activities on improving the amplitude of the beatnote will have to be concluded. 


Images attached to this report
Search Help